uzsbgck:

9474 millimeters to centimeters ( use scienfic notation to answer )

jhonyy9:

use the rule of 3 simple 1 cm = 10 mm x cm = 9474 mm x = 9474*1/10 = 9474/10 = ? cm after you get this so you need to know what is the form of scientific notation for example then you have to write in scientific notation the 376,2 so will be 3,7*10^2

uzsbgck:

9474/ 10 = 947.4?

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @uzsbgck 9474/ 10 = 947.4? \(\color{#0cbb34}{\text{End of Quote}}\) no this is not the form of scientific notation use this my above wrote form of

jhonyy9:

@XioGonz ok ?

XioGonz:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 use the rule of 3 simple 1 cm = 10 mm x cm = 9474 mm x = 9474*1/10 = 9474/10 = ? cm after you get this so you need to know what is the form of scientific notation for example then you have to write in scientific notation the 376,2 so will be 3,7*10^2 \(\color{#0cbb34}{\text{End of Quote}}\) This is correct. I agree!

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @XioGonz \(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 use the rule of 3 simple 1 cm = 10 mm x cm = 9474 mm x = 9474*1/10 = 9474/10 = ? cm after you get this so you need to know what is the form of scientific notation for example then you have to write in scientific notation the 376,2 so will be 3,7*10^2 \(\color{#0cbb34}{\text{End of Quote}}\) This is correct. I agree! \(\color{#0cbb34}{\text{End of Quote}}\) ty so much -

XioGonz:

Npnp! (: